40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Entry  Wed Jul 7 13:05:07 2021, Paco, DailyProgress, 1418 nm AUX ECDL, New aspheric flexures 
    Reply  Fri Jul 9 11:40:42 2021, Paco, DailyProgress, 1418 nm AUX ECDL, New aspheric flexures IMG_1773.jpeg
       Reply  Wed Jul 14 12:21:08 2021, Paco, DailyProgress, 1418 nm AUX ECDL, ECDL lases... and MZ locked mid-fringe selfhomodyne_1418.jpg
          Reply  Wed Aug 4 11:36:30 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL with 2um AOM tests rf_setup.jpgdiffraction_levels.png
             Reply  Tue Aug 10 09:51:44 2021, Paco, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz rf_setup.jpg
                Reply  Thu Aug 12 11:49:59 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz zeroth_order_contour.pdffirst_order_contour.pdf
                   Reply  Tue Aug 17 11:09:29 2021, rana, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz 
                   Reply  Wed Sep 1 13:12:02 2021, Radhika, DailyProgress, 1418 nm AUX ECDL, 1419 nm ECDL AOM diffraction at 95 MHz 
Message ID: 1917     Entry time: Wed Aug 4 11:36:30 2021     In reply to: 1916     Reply to this: 1918
Author: Radhika 
Type: DailyProgress 
Category: 1418 nm AUX ECDL 
Subject: 1419 nm ECDL with 2um AOM tests 

[Radhika, Paco]

In order to transition the ECDL laser noise characterization to a heterodyne setup, we needed to test the AOM (acousto-optic modulator). We wanted to drive the AOM at 80MHz using the Marconi signal generator. Since the AOM has a max driving power of 600 mW, we determined that if we run the Marconi at max output power (13dBm), we saturate the AOM through a variable attenuator and a 5W amplifier. The detailed setup is in Attachment 1.

As we scanned the AOM RF input power, we monitored the mean of the 0th and 1st order power outputs using 2 amplified photodiodes on the scope. Attachment 2 plots the results of the scan; although we noticed the 0th order dropping, we did not see evidence of diffraction in the 1st order. Our suspected theory is that the lost power from the 0th order is due to thermally-driven attenuation inside the AOM (we do not know what is inside the AOM, so this is purely speculative). The next thing we want to try is to add a DC power level to the AOM RF input, but we will double check with Aidan.

Attachment 1: rf_setup.jpg  53 kB  Uploaded Tue Aug 10 11:53:35 2021  | Hide | Hide all
rf_setup.jpg
Attachment 2: diffraction_levels.png  18 kB  Uploaded Tue Aug 10 11:54:27 2021  | Hide | Hide all
diffraction_levels.png
ELOG V3.1.3-